OpenStudy (anonymous):

What is y=x^2/16+1/32x+1? How do I solve this using the Quadratic formula?

OpenStudy (anonymous):

This is a rational equation since it involves fractions to simplify, we have to make all coeficients integers We do this by multiplying all terms by 32 Hence, 32y = 2x2 + x + 32 If you are just trying to identify the equation, it is a RATIONAL EQUATION

OpenStudy (debbieg):

if the equation is (as I think it is)\[y=\frac{ x^2 }{16 }+\frac{ 1 }{ 32 }x+1\]then it is NOT a rational equation, it is a polynomial with rational coefficients. a=1/16, b=1/32, c=1 Plug into quadratic formula. You can't "make all the coefficients integers" because then you have a coefficient other than 1 on y, so can't use the quadratic formula to solve.